Molar Volume of a Gas Room temperature (T2): Barometric pressure: C be & torr Tria .0544 Data Trial 2 Trial 1 Mass of Mg striNeed help solving these last 7 questions. Please use the data above to calculate the answers.
Answer #1

Trial 1 :

Partial pressure of hydrogen gas at room temperature = barometric pressure - vapor pressure = 760.8 - 14.2 torr = 746.6 torr = Partial pressure of hydrogen gas at room temperature in atm , P2 = 746.6 torr = (1 torr = 0.001316 atm) = 0.982 atm

Pstp = 1 atm, Tstp = 273 K , Vstp = V

T1 = 22.3 + 273 = 295.3 K

V1 = 0.0795 lts

Pstp Vstp/Tstp = P2V2/T2

1 atm x V /273 = 0.982 atm x 0.0795 lts/295.3 K

V = 21.312/295.3 = 0.0722lts

Mg(s) + 2HCl(aq) ---> MgCl2(aq) + H2(g)

Moles of Magnesium = mass/molar mass = 0.078 g/24.31 g/mol = 0.0032 moles

Molar ratio of Mg and Hydrogen = 1:1

Moles of hydrogen = 0.0032 moles

Molar volume of hydrogen at STP = V/n = 0.0722 lts/0.0032 mol = 22.56 l/mol

Trial 2 :

Partial pressure of hydrogen gas at room temperature = barometric pressure - vapor pressure = 760.8 - 18.7 torr = 742.1 torr = Partial pressure of hydrogen gas at room temperature in atm , P2 = 742.1 torr = (1 torr = 0.001316 atm) = 0.976 atm

Pstp = 1 atm, Tstp = 273 K , Vstp = V

T1 = 22.3 + 273 = 295.3 K

V1 = 0.0541 lts

Pstp Vstp/Tstp = P2V2/T2

1 atm x V /273 = 0.976 atm x 0.0541 lts/295.3

V = 14.415/295.3 = 0.0488 lts

Mg(s) + 2HCl(aq) ---> MgCl2(aq) + H2(g)

Moles of Magnesium = mass/molar mass = 0.065 g/24.31 g/mol = 0.0027 moles

Molar ratio of Mg and Hydrogen = 1:1

Moles of hydrogen = 0.0027 moles

Molar volume of hydrogen at STP = V/n = 0.0488 lts/0.0027 mol = 18.07 l/mol

Trial 3 :

Partial pressure of hydrogen gas at room temperature = barometric pressure - vapor pressure = 760.8 - 18.7 torr = 742.1 torr = Partial pressure of hydrogen gas at room temperature in atm , P2 = 742.1 torr = (1 torr = 0.001316 atm) = 0.976 atm

Pstp = 1 atm, Tstp = 273 K , Vstp = V

T1 = 22.3 + 273 = 295.3 K

V1 = 0.0564 lts

Pstp Vstp/Tstp = P2V2/T2

1 atm x V /273 = 0.976 atm x 0.0564 lts/295.3

V = 15.028/295.3 = 0.0509 lts

Mg(s) + 2HCl(aq) ---> MgCl2(aq) + H2(g)

Moles of Magnesium = mass/molar mass = 0.0599 g/24.31 g/mol = 0.0025 moles

Molar ratio of Mg and Hydrogen = 1:1

Moles of hydrogen = 0.0025 moles

Molar volume of hydrogen at STP = V/n = 0.0509 lts/0.0025 mol = 20.36 l/mol

Average molar volume = 22.56 + 18.07 + 20.36)/3 = 60.99/3 = 20.33 l/mol

Theoretical value of molar volume of ideal gas at STP = 22.4 l/mol

Percent error = ([exp value - theoretical value])/theoretical value) x 100 = ([20.33-22.4)/22.4) x 100 = 9.24 %